OpenStudy (anonymous):

Does anybody know the value of log3(27)

OpenStudy (mathstudent55):

Do you know the definition of log?

OpenStudy (anonymous):

oh uh no

OpenStudy (mathstudent55):

Here it is: \( \log_b x = y \iff b^y = x\)

OpenStudy (mathmale):

More important is how to obtain that value. One way in which you could do this, probably the easiest way, would be to re-write that 27 as 3^3. Then use the property that y=log3 x and y=3^x are INVERSE FUNCTIONS.

OpenStudy (mathstudent55):

Now using the definition of log, use the values you have for b, x, and y to change the log you are given to the exponential expression to the right.

OpenStudy (mathstudent55):

\(\log_b x = y \iff b^y = x\) \(\log_3 27 = y \iff 3^y = 27\) 3 to what power equals 27?

OpenStudy (anonymous):

oh wow, ok thanks

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

\[\log_{3} 27=\log_{3} 3^{3}=3\log_{3} 3=3*1=3\] \[\log_{a} a=1\]